OpenStudy (anonymous):

Given that P = (-4, 11) and Q = (-5, 8), find the component form and magnitude of vector QP .

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

QP=<-4-(-5), 11-8>=<1,3> magnitude QP=|QP|=sqrt(1^2+3^2)=sqrt(10)